# Long-time asymptotics of (1,3)-sign solitary waves for the damped nonlinear Klein-Gordon equation

### v1

@@ 4.1. Basic properties of $(1,3)$-sign solitary waves
This follows easily from a direct computation using the law of cosines. ∎
@@ 4.2. Distance analysis I
In Subsections 4.2–4.4, we show that no $(1,3)$-sign solitary waves can arise from a configuration of centers $z_{0},z_{1},z_{2},z_{3}$ for which $D\gg 1$ and $\tilde{D}-D\ll D$. We begin with the regime in which $z_{1},z_{2},z_{3}$ form an almost equilateral triangle (in particular, $\tilde{D}-D\ll 1$), and prove that such a configuration cannot yield a $(1,3)$-sign solitary waves. To this end, we first establish the following lemma.
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Furthermore, we assume that there exist $1\ll T_{1}<T_{2}\leq\infty$ such that for $T_{1}\leq t<T_{2}$
$\tilde{D}(t)-\hat{D}(t)<D^{\frac{99}{100}}(t),$ (4.15)
$\max{(\rho_{12}(t),\rho_{13}(t),\rho_{23}(t))}<\tilde{D}(t)+D^{\frac{99}{100}}(t).$ (4.16)
@@
First, since $\rho_{12},\ \rho_{13},\ \rho_{23}$ satisfy (4.16), by Lemma 4.3, we have
$\min{(u_{12}\cdot u_{13},u_{21}\cdot u_{23},u_{31}\cdot u_{32})}-\frac{1}{2}\gtrsim-D^{-\frac{1}{100}}.$
Furthermore, by (4.15) and $\rho_{1},\rho_{2},\rho_{3}\geq\hat{D}$, $c_{12},\ c_{13},\ c_{23}$ satisfy
$\min{(c_{12},c_{13},c_{23})}-\frac{1}{2}\gtrsim-D^{-\frac{1}{100}}.$ (4.18)
@@
Here, we assume $\rho_{1}=\hat{D}$. Then, we have
$|\rho_{1}-D|$ $\lesssim D^{\frac{99}{100}},$
$|\rho_{12}-D|$ $\lesssim D^{\frac{99}{100}},$
$\rho_{2}-D$ $\gtrsim-D^{\frac{99}{100}}.$
@@
$u_{1}\cdot u_{12}+\frac{1}{2}\gtrsim-D^{-\frac{1}{100}}.$ (4.20)
@@
$u_{1}\cdot u_{13}+\frac{1}{2}\gtrsim-D^{-\frac{1}{100}}.$ (4.21)
@@
$|D(t)-D(t^{\prime})|\lesssim D^{\frac{7}{10}}(t^{\prime}),$ (4.27)
@@
$\int_{t^{\prime}}^{T_{3}}\mathcal{F}(D(s))ds\sim D^{\frac{7}{10}}(t^{\prime}).$ (4.28)
Then, we have for $t^{\prime}\leq t\leq T_{3}$
$\max{(\rho_{12}(t),\rho_{13}(t),\rho_{23}(t))}-\tilde{D}(t)\lesssim D^{\frac{7}{10}}(t).$
@@
$\geq\frac{D^{\frac{2}{3}}(t^{\prime})}{3}-D^{\frac{1}{2}}(t^{\prime})\geq 0,$
@@
Furthermore, we assume that $T_{4}<T_{3}$. Then we have
@@
Then, by (4.23) and $|\frac{d}{dt}D^{\frac{2}{3}}|\lesssim D^{-\frac{1}{3}}\mathcal{F}(D)$, we obtain
@@
Therefore, we have $T_{4}=T_{3}$. Then, by Lemma 4.4, we have
$\quad\left(\hat{D}(T_{4})-\frac{\rho_{12}(T_{4})+\rho_{13}(T_{4})+\rho_{23}(T_{4})}{3}\right)-\left(\hat{D}(t^{\prime})-\frac{\rho_{12}(t^{\prime})+\rho_{13}(t^{\prime})+\rho_{23}(t^{\prime})}{3}\right)$
$\gtrsim\int_{t^{\prime}}^{T_{4}}\mathcal{F}(D(s))ds$
$\sim D^{\frac{7}{10}}.$
On the other hand, since we have (4.24) and
$\frac{\rho_{12}(T_{4})+\rho_{13}(T_{4})+\rho_{23}(T_{4})}{3}$ $\geq\tilde{D}(T_{4}),$
@@
$\quad\hat{D}(T_{4})-\tilde{D}(T_{4})+D^{\frac{1}{2}}(t^{\prime})$
$\geq\hat{D}(T_{4})-\tilde{D}(T_{4})+\left(\tilde{D}(t^{\prime})-\hat{D}(t^{\prime})\right)$
$\geq\left(\hat{D}(T_{4})-\frac{\rho_{12}(T_{4})+\rho_{13}(T_{4})+\rho_{23}(T_{4})}{3}\right)-\left(\hat{D}(t^{\prime})-\frac{\rho_{12}(t^{\prime})+\rho_{13}(t^{\prime})+\rho_{23}(t^{\prime})}{3}\right)$
$\quad-D^{\frac{2}{3}}(t^{\prime}).$
Therefore we have
$\tilde{D}(T_{4})-\hat{D}(T_{4})\lesssim-D^{\frac{7}{10}},$
which contradicts to (4.7). Therefore we complete the proof.
@@ 4.3. Distance analysis II
Let $\vec{u}$ be a (1,3)-sign solitary waves . Furthermore, we assume that there exist $1\ll T_{1}<T_{2}\leq\infty$ such that for $T_{1}\leq t<T_{2}$
@@
By symmetry, we may assume that
@@
In this case, by Lemma 4.5, we have $\rho_{23}(t)-D(t)\gtrsim D^{\frac{2}{3}}(t)$ for $T_{1}\leq t\leq T_{2}$. Moreover, since $\rho_{12},\ \rho_{23},\ \rho_{13}$ are continuous and (4.31) holds on $T_{1},T_{2})$, it follows that $\rho_{23}$ is always the maximum of the three. Then, by (3.11) and (3.13), we have
@@
$\quad+\mathcal{F}(\rho_{31})u_{31}\cdot u_{3}+o\left(\mathcal{F}(D)\right).$
@@
$\begin{aligned} \frac{\dot{\rho}_{12}+\dot{\rho}_{13}}{2}&\leq-\frac{\mathcal{F}(\rho_{12})+\mathcal{F}(\rho_{13})}{2}+\frac{u_{1}\cdot u_{21}+u_{1}\cdot u_{31}}{2}\mathcal{F}(\rho_{1})\\ &\quad+\frac{u_{12}\cdot u_{2}}{2}\mathcal{F}(\rho_{2})+\frac{u_{13}\cdot u_{3}}{2}\mathcal{F}(\rho_{3})+\frac{1}{100}\mathcal{F}(D).\end{aligned}$ (4.33)
By (4.31), we have
$|\rho_{12}-D|$ $\lesssim D^{\frac{99}{100}},$
$|\rho_{13}-D|$ $\lesssim D^{\frac{99}{100}},$
$\rho_{1}-D$ $\gtrsim-D^{\frac{99}{100}},$
$\rho_{2}-D$ $\gtrsim-D^{\frac{99}{100}},$
$\rho_{3}-D$ $\gtrsim-D^{\frac{99}{100}}.$
@@
$c_{12}-\frac{1}{2}\gtrsim-D^{-\frac{1}{100}},\ c_{13}-\frac{1}{2}\gtrsim-D^{-\frac{1}{100}}.$ (4.34)
@@
$|\rho_{12}-D|\lesssim D^{\frac{99}{100}},$
$|\rho_{13}-D|\lesssim D^{\frac{99}{100}},$
$|\rho_{1}-D|\lesssim D^{\frac{99}{100}},$
@@
$u_{2}\cdot u_{12}-\frac{1}{2}\gtrsim-D^{-\frac{1}{100}},\ u_{1}\cdot u_{12}+\frac{1}{2}\gtrsim-D^{-\frac{1}{100}},$
$u_{3}\cdot u_{13}-\frac{1}{2}\gtrsim-D^{-\frac{1}{100}},\ u_{1}\cdot u_{13}+\frac{1}{2}\gtrsim-D^{-\frac{1}{100}}.$
@@
Next, we assume $\rho_{2}=\hat{D}$. We have
@@
Furthermore, since $\rho_{2}$ satisfies $|\rho_{2}-D|\lesssim D^{\frac{99}{100}}$, we obtain
@@
Next, we estimate $\frac{\dot{\rho}_{12}+\dot{\rho}_{13}}{2}$. We first treat the terms involving $\rho_{1}$. If $\rho_{1}-D>D^{\frac{99}{100}}$, then $\mathcal{F}(\rho_{1})\lesssim D^{-\frac{1}{100}}\mathcal{F}(D)$; hence, together with (4.39) we have
@@
On the other hand, if $\rho_{1}-D\leq D^{\frac{99}{100}}$, then by Lemma 4.3 we have
@@
Thus, in either case, combining with (4.40) yields
@@
If $\rho_{3}$ attains the minimum, the same argument as in the case where $\rho_{2}$ attains the minimum gives
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Furthermore we have for $T_{1}\leq t\leq T_{2}$
@@
$\mbox{{mid}}{(\rho_{12},\rho_{23},\rho_{13})}\geq\tilde{D}+D^{\frac{1}{3}}.$ (4.43)
@@
In this case, by symmetry and Lemma 4.5, we may assume that $\rho_{23}>\tilde{D}+D^{\frac{2}{3}}$ holds for $T_{1}\leq t\leq T_{2}$. Then, we introduce the following bootstrap estimate
@@
$|D(t)-D(t^{\prime})|\lesssim D^{\frac{2}{5}}(t^{\prime}),$
@@
$\int_{t^{\prime}}^{T_{3}}\mathcal{F}(D(s))ds\sim D^{\frac{2}{5}}(t^{\prime}).$
Then, we have for $t^{\prime}\leq t\leq T_{3}$
$\tilde{D}(t)-\hat{D}(t)\lesssim D^{\frac{2}{5}}(t),$
$\mbox{{mid}}{(\rho_{12}(t),\rho_{13}(t),\rho_{23}(t))}-\tilde{D}(t)\lesssim D^{\frac{2}{5}}(t).$
@@
$\geq\frac{D^{\frac{1}{3}}(t^{\prime})}{2}-D^{\frac{1}{4}}(t^{\prime})\geq 0,$
@@
Furthermore, we assume that $T_{4}<T_{3}$. Then we have
@@
Then, by (4.41) and $|\frac{d}{dt}D^{\frac{1}{3}}|\lesssim D^{-\frac{2}{3}}\mathcal{F}(D)$, we obtain
@@
Therefore, we have $T_{4}=T_{3}$. Then, by Lemma 4.6, we have
$\quad\left(\hat{D}(T_{4})-\frac{\rho_{12}(T_{4})+\rho_{13}(T_{4})}{2}\right)-\left(\hat{D}(t^{\prime})-\frac{\rho_{12}(t^{\prime})+\rho_{13}(t^{\prime})}{2}\right)$
$\gtrsim\int_{t^{\prime}}^{T_{4}}\mathcal{F}(D(s))ds$
$\sim D^{\frac{2}{5}}.$
Furthermore, by (4.42) and (4.43) and
$\frac{\rho_{12}(T_{4})+\rho_{13}(T_{4})}{2}\geq\tilde{D}(T_{4}),$
@@
$\quad\hat{D}(T_{4})-\tilde{D}(T_{4})+D^{\frac{1}{4}}(t^{\prime})$
$\geq\hat{D}(T_{4})-\tilde{D}(T_{4})+\left(\tilde{D}(t^{\prime})-\hat{D}(t^{\prime})\right)$
$\geq\left(\hat{D}(T_{4})-\frac{\rho_{12}(T_{4})+\rho_{13}(T_{4})}{2}\right)-\left(\hat{D}(t^{\prime})-\frac{\rho_{12}(t^{\prime})+\rho_{13}(t^{\prime})}{2}\right)-D^{\frac{1}{3}}(t^{\prime}).$
@@
$\tilde{D}(T_{4})-\hat{D}(T_{4})$ $\lesssim-D^{\frac{2}{5}},$

### v2

@@ 4.1. Basic properties of $(1,3)$-sign solitary waves
We define
$\Lambda_{1}=|\varsigma_{3}-\varsigma_{2}|,\ \Lambda_{2}=|\varsigma_{3}-\varsigma_{1}|,\ \Lambda_{3}=|\varsigma_{2}-\varsigma_{1}|.$
First, we prove (1). By assumption, we have
$\frac{\varsigma_{2}-\varsigma_{1}}{|\varsigma_{2}-\varsigma_{1}|}\cdot\frac{\varsigma_{3}-\varsigma_{1}}{|\varsigma_{3}-\varsigma_{1}|}$ $=\frac{\Lambda_{2}^{2}+\Lambda_{3}^{2}-\Lambda_{1}^{2}}{2\Lambda_{2}\Lambda_{3}}$
$\leq\frac{2(M+M^{\frac{99}{100}})^{2}-(M-M^{\frac{99}{100}})^{2}}{2(M-M^{\frac{99}{100}})^{2}}$
$\leq\frac{M^{2}+7M^{\frac{199}{100}}}{2(M-M^{\frac{99}{100}})^{2}}$
$=\frac{1+\frac{7}{M^{\frac{1}{100}}}}{2\left(1-\frac{1}{M^{\frac{1}{100}}}\right)^{2}}$
$\leq\frac{1}{2}+\frac{5}{M^{\frac{1}{100}}}.$
Furthermore, we have
$\frac{\varsigma_{1}-\varsigma_{2}}{|\varsigma_{1}-\varsigma_{2}|}\cdot\frac{\varsigma_{3}-\varsigma_{2}}{|\varsigma_{3}-\varsigma_{2}|}$ $=\frac{\Lambda_{1}^{2}+\Lambda_{3}^{2}-\Lambda_{2}^{2}}{2\Lambda_{1}\Lambda_{3}}$
$\geq\frac{\Lambda_{1}^{2}+(M-M^{\frac{99}{100}})^{2}-(M+M^{\frac{99}{100}})^{2}}{2\Lambda_{1}\Lambda_{3}}$
$=\frac{\Lambda_{1}}{2\Lambda_{3}}-\frac{4M^{\frac{199}{100}}}{2\Lambda_{1}\Lambda_{3}}$
$\geq\frac{M-M^{\frac{99}{100}}}{2(M+M^{\frac{99}{100}})}-\frac{4M^{\frac{199}{100}}}{2(M-M^{\frac{99}{100}})^{2}}$
$\geq\frac{1}{2}-\frac{5}{M^{\frac{1}{100}}}.$
By the same calculation, we obtain
$\frac{\varsigma_{1}-\varsigma_{3}}{|\varsigma_{1}-\varsigma_{3}|}\cdot\frac{\varsigma_{2}-\varsigma_{3}}{|\varsigma_{2}-\varsigma_{3}|}\geq\frac{1}{2}-\frac{5}{M^{\frac{1}{100}}},$
and therefore we prove (1).
Next we prove (2). By the direct computation, we have
$\frac{\varsigma_{1}-\varsigma_{3}}{|\varsigma_{1}-\varsigma_{3}|}\cdot\frac{\varsigma_{2}-\varsigma_{3}}{|\varsigma_{2}-\varsigma_{3}|}$ $=\frac{\Lambda_{1}^{2}+\Lambda_{2}^{2}-\Lambda_{3}^{2}}{2\Lambda_{1}\Lambda_{2}}$
$\geq\frac{2\Lambda_{1}\Lambda_{2}-(M+M^{\frac{99}{100}})^{2}}{2\Lambda_{1}\Lambda_{2}}$
$\geq 1-\frac{(M+M^{\frac{99}{100}})^{2}}{2(M-M^{\frac{99}{100}})^{2}}$
$=1-\frac{(1+\frac{1}{M^{\frac{1}{100}}})^{2}}{2(1-\frac{1}{M^{\frac{1}{100}}})^{2}}$
$\geq\frac{1}{2}-\frac{5}{M^{\frac{1}{100}}}.$
Therefore we prove (2). ∎
@@ 4.2. Distance analysis I
[…]
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Furthermore, we assume that there exist $1\ll T_{1}<T_{2}\leq\infty$ such that we have for $T_{1}\leq t<T_{2}$
$\tilde{D}(t)-\hat{D}(t)<D^{\frac{9}{10}}(t),$ (4.12)
$\max{(\rho_{12}(t),\rho_{13}(t),\rho_{23}(t))}<\tilde{D}(t)+D^{\frac{9}{10}}(t).$ (4.13)
@@
First, by (4.6) and (4.12), we have for $T_{1}\leq t<T_{2}$
$\left|\tilde{D}(t)-D(t)\right|+\left|\hat{D}(t)-D(t)\right|<D^{\frac{19}{20}}(t).$ (4.15)
Since $\rho_{12},\ \rho_{13},\ \rho_{23}$ satisfy (4.13), by Lemma 4.3 and (4.15), we have
$\min{(u_{12}\cdot u_{13},u_{21}\cdot u_{23},u_{31}\cdot u_{32})}-\frac{1}{2}\geq-10D^{-\frac{1}{100}}.$
Furthermore, by (4.15) and $\rho_{1},\rho_{2},\rho_{3}\geq\hat{D}$, we have
$\min{(\rho_{1},\rho_{2},\rho_{3})}-D\geq-D^{\frac{99}{100}}.$
Therefore, by Lemma 4.3, $c_{12},\ c_{13},\ c_{23}$ satisfy
$\min{(c_{12},c_{13},c_{23})}-\frac{1}{2}\geq-10D^{-\frac{1}{100}}.$ (4.16)
@@
Here, we assume $\rho_{1}=\hat{D}$. Then, by (4.13) and (4.15), we have
$|\rho_{1}-D|$ $\leq D^{\frac{99}{100}},$
$|\rho_{12}-D|$ $\leq D^{\frac{99}{100}},$
$\rho_{2}-D$ $\geq-D^{\frac{99}{100}}.$
@@
$u_{1}\cdot u_{12}+\frac{1}{2}\geq-5D^{-\frac{1}{100}}.$ (4.18)
@@
$u_{1}\cdot u_{13}+\frac{1}{2}\geq-5D^{-\frac{1}{100}}.$ (4.19)
@@
$|D(t)-D(t^{\prime})|\leq 2D^{\frac{7}{10}}(t^{\prime}),$ (4.25)
@@
$c_{\bullet}D^{\frac{7}{10}}(t^{\prime})\leq\int_{t^{\prime}}^{T_{3}}\mathcal{F}(D(s))ds\leq\frac{1}{c_{\bullet}}D^{\frac{7}{10}}(t^{\prime}).$ (4.26)
Furthermore, by Lemma 3.3, we have
$\begin{aligned} \left|\max{(\rho_{12}(t),\rho_{13}(t),\rho_{23}(t))}-\max{(\rho_{12}(t^{\prime}),\rho_{13}(t^{\prime}),\rho_{23}(t^{\prime}))}\right|&\lesssim\int_{t^{\prime}}^{t}\mathcal{F}(D(s))ds,\\ \left|\tilde{D}(t)-\tilde{D}(t^{\prime})\right|&\lesssim\int_{t^{\prime}}^{t}\mathcal{F}(D(s))ds.\end{aligned}$ (4.27)
By (4.6), (4.25), (4.26), and (4.27), we have for $t^{\prime}\leq t\leq T_{3}$
$\tilde{D}(t)-\hat{D}(t)$ $\lesssim\tilde{D}(t^{\prime})-\hat{D}(t^{\prime})+D^{\frac{7}{10}}(t^{\prime})$
$\lesssim D^{\frac{7}{10}}(t^{\prime}),$
and
$\quad\max{(\rho_{12}(t),\rho_{13}(t),\rho_{23}(t))}-\tilde{D}(t)$
$\lesssim\max{(\rho_{12}(t^{\prime}),\rho_{13}(t^{\prime}),\rho_{23}(t^{\prime}))}-\tilde{D}(t^{\prime})+D^{\frac{7}{10}}(t^{\prime}).$
$\lesssim D^{\frac{7}{10}}(t^{\prime}).$
In particular, by (4.25) and $D\gg 1$, we have for $t^{\prime}\leq t\leq T_{3}$
$\begin{aligned} \tilde{D}(t)-\hat{D}(t)<D^{\frac{9}{10}}(t),\\ \max{(\rho_{12}(t),\rho_{13}(t),\rho_{23}(t))}<\tilde{D}(t)+D^{\frac{9}{10}}(t).\end{aligned}$ (4.28)
@@
$\geq\frac{D^{\frac{2}{3}}(t^{\prime})}{3}-D^{\frac{1}{2}}(t^{\prime})>0,$
@@
We assume $T_{4}<T_{3}$. Then, we have
@@
Then, by (4.28), we may apply Lemma 4.4. Therefore, by (4.21) and $|\frac{d}{dt}D^{\frac{2}{3}}|\lesssim D^{-\frac{1}{3}}\mathcal{F}(D)$, we obtain
@@
Therefore, we have $T_{4}=T_{3}$. Then, by Lemma 4.4 and (4.26), there exists $c>0$ such that
$\begin{aligned} &\quad\left(\hat{D}(T_{3})-\frac{\rho_{12}(T_{3})+\rho_{13}(T_{3})+\rho_{23}(T_{3})}{3}\right)-\left(\hat{D}(t^{\prime})-\frac{\rho_{12}(t^{\prime})+\rho_{13}(t^{\prime})+\rho_{23}(t^{\prime})}{3}\right)\\ &\geq c\int_{t^{\prime}}^{T_{3}}\mathcal{F}(D(s))ds\\ &\geq cc_{\bullet}D^{\frac{7}{10}}(t^{\prime}).\end{aligned}$ (4.30)
On the other hand, since we have (4.22), (4.30),
$\frac{\rho_{12}(T_{3})+\rho_{13}(T_{3})+\rho_{23}(T_{3})}{3}$ $\geq\tilde{D}(T_{3}),$
and
@@
$\quad\hat{D}(T_{3})-\tilde{D}(T_{3})+D^{\frac{1}{2}}(t^{\prime})$
$\geq\hat{D}(T_{3})-\tilde{D}(T_{3})+\left(\tilde{D}(t^{\prime})-\hat{D}(t^{\prime})\right)$
$\geq\left(\hat{D}(T_{3})-\frac{\rho_{12}(T_{3})+\rho_{13}(T_{3})+\rho_{23}(T_{3})}{3}\right)-\left(\hat{D}(t^{\prime})-\frac{\rho_{12}(t^{\prime})+\rho_{13}(t^{\prime})+\rho_{23}(t^{\prime})}{3}\right)$
$\quad-D^{\frac{2}{3}}(t^{\prime})$
$\geq cc_{\bullet}D^{\frac{7}{10}}(t^{\prime})-D^{\frac{2}{3}}(t^{\prime}).$
Therefore, we obtain
$\tilde{D}(T_{3})-\hat{D}(T_{3})\leq-cc_{\bullet}D^{\frac{7}{10}}(t^{\prime})+D^{\frac{2}{3}}(t^{\prime})+D^{\frac{1}{2}}(t^{\prime}),$
which contradicts (4.6). Therefore we complete the proof.
@@ 4.3. Distance analysis II
Let $\vec{u}$ be a (1,3)-sign solitary waves . Furthermore, we assume that there exist $1\ll T_{1}<T_{2}\leq\infty$ such that we have for $T_{1}\leq t<T_{2}$
@@
By symmetry, we may assume that we have for $T_{1}\leq t<T_{2}$
@@
In this case, by Lemma 4.5, we have $\rho_{23}(t)-D(t)\gtrsim D^{\frac{2}{3}}(t)$ for $T_{1}\leq t\leq T_{2}$. Moreover, since $\rho_{12},\ \rho_{23},\ \rho_{13}$ are continuous and (4.32) holds on $T_{1},T_{2})$, we may assume that we have for $T_{1}\leq t<T_{2}$
$\rho_{23}(t)=\max{(\rho_{12}(t),\rho_{13}(t),\rho_{23}(t))}.$
Then, by (3.11) and (3.13), we have
@@
$\quad+\mathcal{F}(\rho_{31})u_{31}\cdot u_{3}+o\left(\mathcal{F}(D)\right),$
@@
$\begin{aligned} \frac{\dot{\rho}_{12}+\dot{\rho}_{13}}{2}&\leq-\frac{\mathcal{F}(\rho_{12})+\mathcal{F}(\rho_{13})}{2}+\frac{u_{1}\cdot u_{21}+u_{1}\cdot u_{31}}{2}\mathcal{F}(\rho_{1})\\ &\quad+\frac{u_{12}\cdot u_{2}}{2}\mathcal{F}(\rho_{2})+\frac{u_{13}\cdot u_{3}}{2}\mathcal{F}(\rho_{3})+o\left(\mathcal{F}(D)\right).\end{aligned}$ (4.34)
By (4.6), (4.31), and (4.32), we have
$|\rho_{12}-D|$ $\leq D^{\frac{99}{100}},$
$|\rho_{13}-D|$ $\leq D^{\frac{99}{100}},$
$\rho_{1}-D$ $\geq-D^{\frac{99}{100}},$
$\rho_{2}-D$ $\geq-D^{\frac{99}{100}},$
$\rho_{3}-D$ $\geq-D^{\frac{99}{100}}.$
@@
$c_{12}-\frac{1}{2}\geq-5D^{-\frac{1}{100}},\ c_{13}-\frac{1}{2}\geq-5D^{-\frac{1}{100}}.$ (4.35)
@@
$|\rho_{12}-D|\leq D^{\frac{99}{100}},$
$|\rho_{13}-D|\leq D^{\frac{99}{100}},$
$|\rho_{1}-D|\leq D^{\frac{99}{100}},$
@@
$u_{2}\cdot u_{12}-\frac{1}{2}\geq-5D^{-\frac{1}{100}},\ u_{1}\cdot u_{12}+\frac{1}{2}\geq-5D^{-\frac{1}{100}},$
$u_{3}\cdot u_{13}-\frac{1}{2}\geq-5D^{-\frac{1}{100}},\ u_{1}\cdot u_{13}+\frac{1}{2}\geq-5D^{-\frac{1}{100}}.$
@@
Next, suppose that $\rho_{2}=\hat{D}$. We have
@@
Furthermore, since $\rho_{2}$ satisfies $|\rho_{2}-D|\leq D^{\frac{99}{100}}$, we obtain
@@
Next, we estimate $\frac{\dot{\rho}_{12}+\dot{\rho}_{13}}{2}$. We first treat the terms involving $\rho_{1}$. If $\rho_{1}$ satisfies $\rho_{1}-D>D^{\frac{99}{100}}$, then we have $\mathcal{F}(\rho_{1})\lesssim D^{-\frac{1}{100}}\mathcal{F}(D)$; hence, together with (4.40) we have
@@
On the other hand, if $\rho_{1}$ satisfies $\rho_{1}-D\leq D^{\frac{99}{100}}$, then by Lemma 4.3 we have
@@
Thus, in either case, combining the above estimate with (4.41) yields
@@
If $\rho_{3}=\hat{D}$, the same argument as in the case where $\rho_{2}$ attains the minimum gives
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Furthermore, we assume that we have for $T_{1}\leq t\leq T_{2}$
@@
$\mbox{{mid}}{(\rho_{12}(t),\rho_{13}(t),\rho_{23}(t))}\geq\tilde{D}(t)+D^{\frac{1}{3}}(t).$ (4.44)
@@
In this case, by symmetry and Lemma 4.5 and continuity, we may assume that $\rho_{23}>\tilde{D}+D^{\frac{2}{3}}$ holds for $T_{1}\leq t\leq T_{2}$. Then, we introduce the following bootstrap estimate
@@
$|D(t)-D(t^{\prime})|\leq 2D^{\frac{2}{5}}(t^{\prime}),$
@@
$c_{\bullet}D^{\frac{2}{5}}(t^{\prime})\leq\int_{t^{\prime}}^{T_{3}}\mathcal{F}(D(s))ds\leq\frac{1}{c_{\bullet}}D^{\frac{2}{5}}(t^{\prime}).$
Then, by Lemma 3.3, (4.6), (4.43), and (4.45), we have for $t^{\prime}\leq t\leq T_{3}$
$\tilde{D}(t)-\hat{D}(t)$ $\lesssim\tilde{D}(t^{\prime})-\hat{D}(t^{\prime})+D^{\frac{2}{5}}(t^{\prime})$
$\lesssim D^{\frac{2}{5}}(t),$
$\mbox{{mid}}{(\rho_{12}(t),\rho_{13}(t),\rho_{23}(t))}-\tilde{D}(t)$ $\lesssim\mbox{{mid}}{(\rho_{12}(t^{\prime}),\rho_{13}(t^{\prime}),\rho_{23}(t^{\prime}))}-\tilde{D}(t^{\prime})+D^{\frac{2}{5}}(t^{\prime})$
$\lesssim D^{\frac{2}{5}}(t).$
In particular, we have for $t^{\prime}\leq t\leq T_{3}$
$\begin{aligned} \tilde{D}(t)-\hat{D}(t)&<D^{\frac{1}{2}}(t),\\ \mbox{{mid}}{(\rho_{12}(t),\rho_{13}(t),\rho_{23}(t))}&<\tilde{D}(t)+D^{\frac{1}{2}}(t).\end{aligned}$ (4.47)
@@
$\geq\frac{D^{\frac{1}{3}}(t^{\prime})}{2}-D^{\frac{1}{4}}(t^{\prime})>0,$
@@
We assume that $T_{4}<T_{3}$. Then, we have
@@
Then, by (4.47), we may apply Lemma 4.6. Therefore, by (4.42) and $|\frac{d}{dt}D^{\frac{1}{3}}|\lesssim D^{-\frac{2}{3}}\mathcal{F}(D)$, we obtain
@@
Therefore, we have $T_{4}=T_{3}$. Then, by Lemma 4.6, there exists $c>0$ such that
$\begin{aligned} &\quad\left(\hat{D}(T_{3})-\frac{\rho_{12}(T_{3})+\rho_{13}(T_{3})}{2}\right)-\left(\hat{D}(t^{\prime})-\frac{\rho_{12}(t^{\prime})+\rho_{13}(t^{\prime})}{2}\right)\\ &\geq c\int_{t^{\prime}}^{T_{3}}\mathcal{F}(D(s))ds\\ &\geq cc_{\bullet}D^{\frac{2}{5}}(t^{\prime}).\end{aligned}$ (4.49)
Furthermore, by (4.43), (4.45), (4.49), and
$\frac{\rho_{12}(T_{3})+\rho_{13}(T_{3})}{2}\geq\tilde{D}(T_{3}),$
@@
$\quad\hat{D}(T_{3})-\tilde{D}(T_{3})+D^{\frac{1}{4}}(t^{\prime})$
$\geq\hat{D}(T_{3})-\tilde{D}(T_{3})+\left(\tilde{D}(t^{\prime})-\hat{D}(t^{\prime})\right)$
$\geq\left(\hat{D}(T_{3})-\frac{\rho_{12}(T_{3})+\rho_{13}(T_{3})}{2}\right)-\left(\hat{D}(t^{\prime})-\frac{\rho_{12}(t^{\prime})+\rho_{13}(t^{\prime})}{2}\right)-D^{\frac{1}{3}}(t^{\prime})$
$\geq cc_{\bullet}D^{\frac{2}{5}}(t^{\prime})-D^{\frac{1}{3}}(t^{\prime}).$
@@
$\tilde{D}(T_{3})-\hat{D}(T_{3})$ $\leq-cc_{\bullet}D^{\frac{2}{5}}(t^{\prime})+D^{\frac{1}{3}}(t^{\prime})+D^{\frac{1}{4}}(t^{\prime}),$